Evaluate the integeral
∫((2x³ -4x -8)/ (x⁴ -x³ +4x² -4x)) dx
1
Expert's answer
2020-07-05T17:19:46-0400

"I = \int\frac{2x^3-4x-8}{x^4-x^3+4x^2-4x}dx."


This is a fractional-rational function "\frac{2x^3-4x-8}{x^4-x^3+4x^2-4x}". Factorize the denominator:


"x^4-x^3+4x^2-4x = x^3(x-1)+4x(x-1)="

"=(x-1)(x^3+4x)=x(x-1)(x^2+4)."


So, we have "\frac{2x^3-4x-8}{x^4-x^3+4x^2-4x}=\frac{2x^3-4x-8}{x(x-1)(x^2+4)}".


Let's build a decomposition:


"\frac{2x^3-4x-8}{x(x-1)(x^2+4)}=\frac{A}{x}+\frac{B}{x-1}+\frac{Cx+D}{x^2+4}".

Let's bring the right part to the common denominator:


"\frac{}{}""\frac{A(x-1)(x^2+4)+Bx(x^2+4)+(Cx+D)(x-1)x}{x(x-1)(x^2+4)}".


Open the brackets in the numerator:

"A(x^3+4x-x^2-4)+Bx^3+4Bx+Cx^3-"

"-Cx^2+Dx^2-Dx".


Simplify the expression:

"x^3(A+B+C)+x^2(-A-C+D)+"

"+x(4A+4B-D)+x^0(-4A)" .


So, we have:

"2x^3-4x-8=x^3(A+B+C)+"

"+x^2(-A-C+D)+x(4A+4B-D)+"

"+x^0(-4A)".


Let's collect summands with the same degrees:

"A+B+C=2" ,

"D-A-C=0" ,

"4A+4B-D=-4" ,

"-4A=-8".


From the last equation we find "A = 2".

Now, we have:

"B+C=0",

"D-C=2",

"4B-D=-12".


Further:

"B=-\u0421",

"D=2+C",

"4B-2+B=-12" .


So, from the last equation we find "B=-2". And therefore "C=2, D=4".

Eventually:


"I = \int\frac{2x^3-4x-8}{x^4-x^3+4x^2-4x}dx ="


"=\int\frac{A}{x}+\frac{B}{x-1}+\frac{Cx+D}{x^2+4}=\int\frac{2}{x}dx+\int\frac{-2}{x-1}dx+"


"+\int\frac{2x+4}{x^2+4}dx".


Let's count each integral separately:


"\int\frac{2}{x}dx = 2ln(x) +C_1".


"\int\frac{-2}{x-1}dx = -2ln(x-1)+C_2".


"\int\frac{2x+4}{x^2+4}dx=\int\frac{2x}{x^2+4}dx+4\int\frac{1}{x^2+4}dx=\int\frac{d(x^2+4)}{x^2+4}+"


"+2arctan(\frac{x}{2}) +C_4=ln(x^2+4)+C_3+"


"2arctan(\frac{x}{2})+C_4".


Answer:

"I=2ln(x)-2ln(x-1)+ln(x^2+4)+"

"+2arctan(\frac{x}{2})+C".